OpenStudy (anonymous):

Simplify quantity 4 times x minus 20 all over 4 times x. A:negative 20 over 1 B:quantity x minus 5 all over 4 C:quantity x minus 5 all over x D:quantity x minus 5 all over 4 times x

OpenStudy (anonymous):

\[\huge \color{blue}{\frac{4x-20}{4x} = \frac{\cancel{4} (x - 5)}{\cancel{4}x} = \frac{x-5}{x}}\]

OpenStudy (anonymous):

It is C..

OpenStudy (anonymous):

yeah,,u'r right,,i was wrong by cancelling..sorry